B-9[B-8(4-B)]=608

I'm getting 64.9, I think I'm doing a step wrong?

Answer by nerdybill(7384)

You must apply the "order of operations" (PEMDAS):
B-9[B-8(4-B)]=608
Distribute the -8 to terms inside parenthesis:
B-9[B-32+8B]=608
Combining like-terms inside brackets:
B-9[9B-32]=608
Distribute the -9 to terms inside brackets:
B-81B+288=608
Combining like-terms:
-80B+288=608
Subtracting 288 from both sides
-80B=608-288
-80B=320
Dividing both sides by -80
B = -4